### v2

# Measurement of the $g$ factor of ground-state 87Sr at the parts-per-million level using co-trapped ultracold atoms

Abstract
We demonstrate nuclear magnetic resonance of optically trapped ground-state ultracold 87Sr atoms. Using a scheme in which a cloud of ultracold 87Rb is co-trapped nearby, we improve the determination of the nuclear $g$ factor, $g_{I}$, of atomic 87Sr by more than two orders of magnitude, reaching accuracy at the parts-per-million level. We achieve similar accuracy in the ratio of relevant $g$ factors between Rb and Sr. This establishes ultracold 87Sr as an excellent linear in-vacuum magnetometer. More generally, our work demonstrates how ultracold neutral atoms can be used for the precise determination of highly relevant atomic and nuclear magnetic properties. These results are important for ongoing efforts toward quantum simulation, quantum computation and optical atomic clocks employing 87Sr, and other neutral alkaline-earth and alkaline-earth-like atoms.
Alkaline-earth and alkaline-earth-like atoms are excellent candidates for quantum simulation of many-body physics [1], quantum computing [2] and atom interferometry [3, 4]. They have been pivotal in realizing state-of-the-art optical atomic clocks [5] and molecular clocks [6]. Many of these applications strongly rely on having a singlet electronic ground state and long-lived triplet metastable states. Among these atoms, fermionic 87Sr is especially suited for optical atomic clocks [7, 8, 9, 10], and also stands out because it has the largest nuclear spin quantum number, $I=9/2$. The tenfold degeneracy arising from this large nuclear spin gives rise to SU(10) spin symmetry. 87Sr is thus an ideal platform to study SU($N$) spin Hamiltonians with a great degree of control and isolation—for example implementing Kondo lattice models in an optical lattice [11]. Single 87Sr atoms trapped in optical tweezers hold great potential for realizing arrays of higher-dimensional qubits, i.e. qudits, with up to ten states [12, 13, 14]. A high degree of control in state preparation and manipulation of the nuclear spin states are among the requirements for realizing these systems.
To control and manipulate the nuclear spin states in 87Sr coherently, a key quantity is the splitting of the energy levels in the presence of a magnetic field, the Zeeman effect. Alkaline-earth(-like) neutral atoms experience weak splittings and/or shifts in an external magnetic field due to the absence of electronic (orbital and spin) angular momentum in the (singlet) electronic ground state. The splitting is determined by the nuclear $g$ factor, $g_{I}$. This factor may be written as $g_{I}=\mu_{I}(1-\sigma_{d})/\mu_{B}I$ [15], where $\mu_{B}$ is the Bohr magneton, $\mu_{I}$ is the magnetic moment of the bare nucleus, and $\sigma_{d}$ is the diamagnetic shielding factor due to the surrounding electron cloud. For neutral 87Sr atoms, the value of $g_{I}$ is known experimentally at the $10^{-4}$ level [16, 17].
Figure 1: Schematic of the experimental setup used to measure $g_{I}$ of 87Sr in the electronic ground state. Clouds of Rb and Sr atoms are confined in respective crossed optical dipole traps, where the (near-)vertical dipole trap beam is common to both traps. The microwave radiation is emitted using a dedicated microwave antenna, whereas the radio frequency radiation is emitted using coils carrying alternating current.
Figure 2: (a) Breit-Rabi diagram for 87Rb in the electronic ground state. The states are labeled by their total magnetic quantum number $m$, and (at low field) by their total angular momentum $F$. The violet arrow marks the hyperfine transition between the two m=1m=-1 states used here for (co)magnetometry. (b) Zeeman splitting of 87Sr in the electronic ground state. The violet arrows show the transitions between nuclear spin states, labeled by their nuclear magnetic quantum number $m_{I}$. Note the differences between (a) and (b), particularly regarding energy scale. These are due to the total electron spin being zero in Sr, while it is nonzero in Rb.
Precision measurement of the $g$ factor of atoms and ions can potentially test quantum electrodynamics (QED) corrections to atomic Hamiltonians involving magnetic interactions. Such measurements are being pursued as a tool for searching for new physics beyond the standard model [18, 19]. By measuring the ratio of the $g$ factor of different isotopes, most of the bound-state QED terms are canceled out, allowing one to probe nuclear effects [20, 21]. The ratio of the microwave clock transitions of 87Rb and 133Cs directly depends on the fine structure constant and the $g$ factor of both atoms and this is used to set limits on the temporal variation of fundamental constants. Improving the accuracy of $g$-factor values enhances the performance of atomic magnetometers, which rely on the measurement of the Zeeman shift to measure the magnetic field. Atomic magnetometers are often used in a comagnetometer configuration together with noble gases or other isotopes, where the samples overlap spatially to gain maximum common-mode sensitivity for the in situ magnetic field along with the measurement of interest [22]. Comagnetometers are used for various fundamental physics experiments, such as probing Lorentz and CPT violations, searching for permanent electric dipole moments, and exotic spin-dependent interactions [23].
Here, we report the observation of nuclear magnetic resonance (NMR) of optically trapped ultracold 87Sr atoms. We have measured NMR transitions at magnetic fields ranging from 100 G to 1050 G, resulting in the variation of transition frequencies from 18 kHz to 194 kHz. In a comagnetometer-like configuration, we use hyperfine transitions in 87Rb to determine the magnetic field in situ, with mG accuracy. Combining these measurements, we determine the $g$ factor of neutral 87Sr in its electronic ground state with unprecedented accuracy, at the parts-per-million (ppm) level.
This improves by two orders of magnitude upon the previous state of the art [24, 25, 26], which stood the test of time for over 50 years. Thus, it provides important input for the detailed characterization of the optical clock transition in 87Sr [27, 15]. More generally, this work shows how ultracold neutral atoms can be used to bring orders of magnitude improvement in the experimental accuracy of highly relevant atomic and nuclear magnetic properties, an area for which typically trapped ions are used [28, 29].
[…]
Figure 3: (a) Spectrum of the 87Rb hyperfine transition from the lower m=1m=-1 state to the upper m=1m=-1 state at an inferred magnetic field of 500.544 G, with Gaussian fit. (b) 87Sr nuclear magnetic resonance spectrum measured simultaneously in a comagnetometer-like configuration. Solid lines are Gaussian fits to the data.
The Zeeman structures of the electronic ground states of 87Rb and 87Sr are illustrated in Fig. 2. In the experiments, the atoms are prepared in a specific $m$ state. The 87Rb atoms have $\sim 90\%$ population in the lower m=1m=-1 state, while for 87Sr, after loading the dipole trap, we optically pump $\sim 95\%$ of the atoms to the $m_{I}=+9/2$ state. We ramp the magnetic field to the desired value and then perform simultaneous spectroscopy by applying radiofrequency (RF) and microwave (MW) frequency fields to both atomic clouds simultaneously, with pulse durations between 1.5 s and 3 s. Because of the vastly different resonant frequencies (see the vertical scales in Fig. 2), 87Sr is only sensitive to the RF, while Rb is only sensitive to the MW field. After the spectroscopy pulses, the magnetic field is ramped down on a timescale of 250 ms; next the atoms are released by switching off the dipole traps. During the subsequent (8 ms) time of flight, we employ an optical levitation technique to achieve spin-resolved detection of 87Sr. Specifically, we selectively levitate the $m_{I}=+9/2$ state using the scattering force of a laser driving the $\sigma^{+}$ $^{1}S_{0}$ – $^{3}P_{1}$ cycling transition [31], while the rest of the spin states fall under gravity. A small horizontal offset between the Rb and Sr traps ensures that the Rb does not fall through the center of the optically levitated Sr cloud, avoiding Rb-Sr collisions.
After the time of flight, the Sr clouds are spatially separated enough to distinguish the $m_{I}=+9/2$ state from the rest of the spin states, and to extract the resulting fraction of Sr atoms in $m_{I}=+9/2$ by absorption imaging with 461-nm light. For Rb, the atoms transferred to the $F=2$ manifold are imaged using the cooling transition of Rb at 780 nm. The cycle time of the experiment is $\sim 40$ s.
A typical Rb magnetometer spectrum and the corresponding NMR spectrum of 87Sr are shown in Figs. 3(a) and (b) respectively. Each data point represents the average of a set of several (typically six) experimental cycles, and error bars represent the standard error of each set. For such spectra, the (near-)resonant RF and MW fields were applied with a square envelope (constant amplitude). We can readily achieve sub-Hz Gaussian linewidths for the Sr NMR transition.
We fit Gaussian line shapes to the Sr and the Rb spectra. From the peak position of the fit to the Rb spectrum we extract the magnitude of the local magnetic field at the position of the atoms, while the width of the Rb spectrum is used to determine an upper bound of the residual magnetic field fluctuations over the measurement duration of 1.5 to 2 hours. The data are acquired in a sequential manner, with six scans from low to high frequency. Using the Breit-Rabi formula, the magnetic-field dependence of the transition frequency $f(^{87}$Rb) between the two m=1m=-1 hyperfine states used here can be written as
$f(^{87}{\rm Rb})=f_{0}(^{87}{\rm Rb})\sqrt{(1-x)^{2}+x},$ (1)
[…]
We perform these measurements at magnetic field settings ranging from 100 G to 1050 G to determine the frequency of the 87Sr NMR transitions as a function of magnetic field. Table 1 in the End Matter summarizes the resonance frequencies and rms widths obtained from Gaussian fits to the experimental spectra of both 87Rb and 87Sr. Taking the magnetic field strengths inferred from the 87Rb data, and using the NMR transition frequency $f(^{87}{\rm Sr})=|g_{I,{\rm Sr}}|\mu_{B}B/h$ between the energy levels
$E_{m_{I}}({\rm Sr})=-m_{I}g_{I,{\rm Sr}}\mu_{B}B,$ (2)
the slope of a linear fit gives us the value of $g_{I}$ of 87Sr. From a linear fit to these data we obtain a value of 184.43332(25) Hz/G, achieving a relative precision of 1.4 ppm. See also Fig. 4 and its discussion in the End Matter.
We now address possible systematic effects in the above determination of the linear Zeeman splitting of ground-state 87Sr. The main systematic shift arises from the residual magnetic field difference between the locations of the Rb and Sr atoms. Because of the large (near-resonant) s-wave scattering length between 87Rb and 87Sr [36, 37, 30], there would be substantial shifts (and three-body losses) if we held the Rb and Sr atoms at the same place at the same time, precluding sufficiently long interrogation times. Instead, as mentioned above, we use a small (sub-mm) displacement of the Rb magnetometer to spatially separate it from the Sr cloud. In this modified comagnetometer arrangement, the frequency measurements are done coincident in time and almost coincident in space. The presence of any stray magnetic field gradient near the atomic cloud thus would lead to a shift in the magnetic field at the position of the Rb cloud compared to the position of the Sr cloud.
To characterize the strength of the relevant magnetic field difference, we performed magnetometer measurements on Rb at different positions using a movable dipole trap (in the absence of the Sr cloud) and a magnetic field setting of 1001 G. The position of maximum spatial overlap of the Rb and Sr clouds was determined by scanning the relative position in a separate experiment, maximizing the strong (resonant three-body) losses that occur when Rb and Sr are in contact. We extract a residual gradient of 56(8) mG/mm and attribute this to a slight inhomogeneity of the magnetic field from the main coils [38] at the location of the atoms. To account for this inhomogeneity, we apply a relative correction of $17(3)\times 10^{-6}$ to $B_{Rb}$ to obtain the magnetic field at the Sr atoms.
Further systematic effects we have considered are differential light shifts and density shifts between the coupled spin states of Rb and similarly for the Sr atoms in our optical dipole traps. These are all found to be negligibly small compared to the one arising from the above magnetic field inhomogeneity. A more detailed discussion can be found in the End Matter, and the results are summarized in Table 2 there.
Using Eq. (2) then leads to
$g_{I,{\rm Sr}}=-131.7712(2)_{\rm stat}(3)_{\rm sys}\times 10^{-6}.$ (3)
Expressed as an effective magnetic moment of an isolated neutral 87Sr atom, this yields $\mu_{\rm eff}=g_{I,{\rm Sr}}\mu_{B}I=-1.088784(3)\mu_{N}$, with $\mu_{N}$ the nuclear magneton. This value can be compared to the existing reference value for neutral Sr atoms, measured using optical pumping [16, 17], namely $\mu_{\rm eff}=-1.08859(65)\mu_{N}$. Our determination thus is consistent with this value, and improves on the relative accuracy of $g_{I}$ and $\mu_{\rm eff}$ of 87Sr by more than two orders of magnitude.
Another interesting reference value is the experimental determination of the NMR frequency of 87Sr ions in aqueous solution, yielding $\mu_{\rm eff}=-1.089274(7)\mu_{N}$ [24, 25]. For a direct comparison to our data, the difference in diamagnetic shielding $\sigma_{d}$ between free atoms and ions in solution would have to be known with sufficient accuracy, posing a challenge to the current state of the art of calculating $\sigma_{d}$. We elaborate on this point in the End Matter. It would also be very interesting to connect our result for 87Sr to the recent experimental and theoretical determination [4] of the diamagnetism of the other (naturally occurring, even-numbered) isotopes of Sr, which do not have nuclear spin.
An alternative approach in analyzing our data is to eliminate the magnetic field $B$ and obtain a direct relation between $f(^{87}$Sr) and $f(^{87}$Rb), by substituting $B$ from Eq. (2), into Eq. (1) (including the above systematic correction to $B$). Fitting the resulting equation to the data leads to a precise determination of the ratio of magnetic moments
$\frac{g_{I,{\rm Sr}}}{g_{J}-g_{I,Rb}}=-65.7762(2)\times 10^{-6},$ (4)
with the uncertainty dominated by our present systematic uncertainty in the numerator. The uncertainty in the denominator (i.e. the current relative uncertainty in the literature value of $g_{J}-g_{I}$ of ground-state 87Rb) is at the $10^{-8}$ level, with similar-sized contributions from both $g_{J}$ and $g_{I}$ (see values quoted above). We expect that, with improvements of our method and with acquiring further statistics, we should be able to reach the $10^{-8}$ accuracy level in both $g_{I,{\rm Sr}}$ and in the ratio of Eq. (4), i.e., similar to the established accuracy of $g_{J}$ values in the alkalis [35].
Future experimental improvements that we envision include bringing the ultracold Rb and Sr closer together ($\approx 0.1$ mm separation) and, more importantly, closer to the midpoint between the two magnetic coils, suppressing the effect of magnetic field inhomogeneity that is now dominating our systematic uncertainty. Ideally, the two clouds would be at equal distance from the midpoint where the magnetic field magnitude $B$ is at its extremum (minimum), so that the magnetic field experienced by the respective clouds would be equal (up to at least third order in spatial derivatives of $B$). Acquiring further statistics, in particular at high magnetic fields where the sensitivity is greatest, should allow us to also significantly reduce the statistical uncertainty in $g_{I,{\rm Sr}}$. Some of the other effects that might limit accuracy and precision in next-generation experiments are identified and analyzed in the End Matter.
Our method of NMR on the ten nuclear spin states of optically trapped 87Sr can be used as a critical tool to prepare coherent superpositions among these states, which otherwise are only accessible via Raman transitions in the optical domain. Further, in the presence of a large tensor light shift, the transition frequency between different nuclear spin states becomes distinct, opening up the possibility of using 87Sr as a qudecimal with far more quantum control [12, 13]. These transitions can also serve as an essential asset to the toolbox for realizing SU($N$) spin Hamiltonians with alkaline-earth atoms.
In addition to the relevance of magnetic properties of neutral atoms such as 87Sr for applications in quantum information science and technology, accurate ratios of $g$ factors such as those of Eq. (4) may also offer a different window of opportunity to search for new physics [18]— for instance involving exotic nuclear-spin-dependent couplings [23]. In this context, we point out that Eq. (4) is a ratio of a (predominantly) nuclear property (of 87Sr) and a (predominantly) electronic one (of 87Rb). This will be a generally interesting feature for constraining possible variations in (ratios of) fundamental constants of nature, especially if we can improve the accuracy and precision of our data further.
The difference in linear Zeeman splitting among the $^{1}S_{0}$ and $^{3}P_{0}$ clock states of 87Sr has previously been determined at the 0.4-Hz/G level [27, 15]. Our determination of the absolute value of $g_{I}$ of $^{1}S_{0}$ (and hence its linear Zeeman splitting) is accurate at the mHz/G level, and thus can be used to further constrain the $g$ factor of the $^{3}P_{0}$ state. Looking ahead, we expect that we can extend our experimental methods to perform a precision measurement of $g_{I}$ of $^{3}P_{0}$ using NMR of ultracold 87Sr. We expect to reach similar level of precision as our present determination for $^{1}S_{0}$. This could involve using NMR of $^{1}S_{0}$ for calibration and accuracy (i.e. eliminating the need for Rb as a co-trapped reference species). A more precise value of the $^{3}P_{0}$ $g$ factor is important for characterizing the effects of Zeeman splittings and shifts at the present and near-future levels of systematic uncertainty for optical atomic clocks based on 87Sr [9].
In conclusion, we have measured the $g_{I}$ factor of neutral 87Sr using NMR with an unprecedented accuracy at the ppm level using a modified comagnetometer scheme. This reduces the uncertainty in this $g_{I}$ by more than two orders of magnitude, and poses a challenge to calculations of the diamagnetic shielding factor $\sigma_{d}$. This determination of $g_{I}$ establishes 87Sr as an excellent linear in-vacuum magnetometer, and will be important for applications in quantum simulation, computing, metrology and sensing with fermionic neutral strontium, and in particular for evaluating the magnetic sensitivity of optical lattice clocks based on these atoms. Our work shows how ultracold neutral atoms can be used for precise and accurate determination of highly relevant atomic and nuclear magnetic data. These methods should be readily applicable to other alkaline-earth and alkaline-earth-like atoms.
Acknowledgments—We thank R.J.C. Spreeuw and the participants of the ICOLS 2025 conference for stimulating discussions. This research was funded by NWO Programme “Atomic quantum simulators 2.0” (Project No. 680.92.18.05) and EU project MosaiQC (Project No. 860579). This work was supported by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. The research of AU is partially funded by NWO Grant No.ÑGF.1623.23.025 (“Qudits in theory and experiment”).